# Soft decision trees for survival analysis

### v1

@@ 5.2 Comparison with three benchmark survival tree methods
In the numerical experiments, we consider SSTs with parametric or spline-based semiparametric models for the survival functions. SSTs involving the parametric Exponential, Weibull, and LogLogistic distributions are denoted as Exp, W, and Llog, respectively. For the semiparametric survival function models, SSTs based on the PO and PH assumptions are referred to as PO and PH, respectively. To show the impact of an appropriate initial solution on the calibration and discrimination measures, we also report the results for the Llog parametric distribution when NODEC-DR-SST is combined with the K-means-based initialization procedure proposed in (Consolo et al., 2025), referred to as Llog-init.
@@
Tables 3 and 4 report the average testing $C_{H}$ and IBS for each method. The arithmetic averages over all datasets appear at the bottom of each table.
Concerning the $C_{H}$ discrimination measure in Table 3, for which higher values correspond to better solutions, Llog and Llog-init (with the clustering-based initialization) achieve the best average testing accuracy among the SSTs with parametric or semiparametric models at leaf nodes, and perform better than the three benchmarking methods. In particular, the average testing values for Llog, CTree (the best benchmarking method), and both the comparable semiparametric models PO and PH are, $C_{H}=0.744$, $C_{H}=0.711$, and $C_{H}=0.735$ and $C_{H}=0.736$. Overall, Llog outperforms CTree on 14 out of the 15 datasets, while PO and PH on 13 of them. The Employee dataset is the only one for which SkSurv, CTree, and RPART yield better $C_{H}$ values than the SST alternatives for all the four performance measures. Note that Llog-init, which includes an appropriate initialization procedure, leads to a one percent improvement with respect to Llog for the three discrimination measures, see B.
@@
Table 3: Testing accuracy in terms of $C_{H}$ discriminant measure (the higher the better). The comparison includes SSTs of depth $D=2$ with parametric distributions (Exp, W, Llog) and spline-based semiparametric models (PO and PH), as well as the three benchmarking survival tree methods (SkSurv, CTree, and RPART) with depth $D=5$. Llog-init refers to Llog with the clustering-based initialization procedure.
As to the calibration testing measure IBS in Table 4, for which lower values correspond to better solutions, the spline-based semiparametric SSTs, PO and PH, outperform the benchmarking methods SkSurv, CTree, and RPART, as well as the parametric SSTs. Among the benchmarking methods, CTree achieves the lowest IBS ($0.11$), comparable to Llog ($0.111$), but it turns out to be less accurate than PO and PH, which yield $\text{IBS}=0.074$ and $\text{IBS}=0.075$, respectively. Notice that the K-means-based initialization procedure further enhances the performance of Llog, achieving an average IBS of $0.107$ compared to $0.111$ with no initialization.
@@
Table 4: Testing accuracy in terms of calibration measure IBS (the lower the better). The comparison includes SSTs of depth $D=2$ with parametric distributions (Exp, W, Llog) and spline-based semiparametric models (PO and PH), as well as the three benchmarking survival tree methods (SkSurv, CTree, and RPART) with depth $D=5$. Llog-init refers to Llog with the clustering-based initialization procedure.
@@ 5.3 Enhancing interpretability
The aim of this section is to demonstrate how the SST model enhances interpretability by analyzing the clusters of survival functions corresponding to each leaf node. Unlike other single-tree methods, such as greedy algorithms (Hothorn et al., 2015) or more recent approaches employing advanced optimization techniques (Bertsimas et al., 2022; Huisman et al., 2024; Zhang et al., 2024), the SST is sufficiently flexible to define distinct survival functions for each data point within a leaf node. This approach enables the exploration of differences both across and within leaf node clusters. For the sake of conciseness, the analysis focuses on the $9$ dataset, with results for other $6$ datasets presented in the E.
@@ 5.4 Fairness
Formulation (4) is optimized using the node-based decomposition training algorithm on a socio-economic dataset where a sensitive group is identifiable. Since the datasets discussed in Section 5.1 involve applications where it is not natural to identify a sensitive group, this section focuses on the Unemployment dataset from Romeo (1999). The Unemployment dataset consists of a four-month panel of revised Current Population Survey data from September 1993, where each observation corresponds to an unemployed individual searching for a job. Each row includes binary variables indicating the reason for unemployment (e.g., whether the individual is a job loser, job leaver, or labor force reentrant), the search method employed, and their demographic status. The dataset consists of $452$ data points, $7$ features, with a censoring level of $56.6\%$. Gender is selected as the sensitive feature for group $S$ [^6] This feature was not considered for SST training and hence, it can be used to predict survival functions of forthcoming individuals without even knowing this feature.
@@
For the sake of illustration, we consider SST with the Llog parametric distribution and a single fold of the 5-fold cross-validation. We also perform a single run without applying the clustering-based initialization procedure. It is important to highlight that any alternative parametric or semiparametric choice of survival function for the leaf nodes is feasible. The decision to present a single run is motivated solely by the aim of clearly illustrating the distribution of clusters across the 4 leaf nodes and its trend as the hyperparameter $\rho$, associated with the fairness penalty, increases. The $\rho$ hyperparameter starts from 00 and reaches $\frac{20}{N_{M}N_{F}}$ in five steps, defined by $\frac{1}{N_{M}N_{F}}[0,1,5,10,15,20]$, where $N_{M}$ and $N_{F}$ correspond to the number of males and females in the training dataset.
@@ 6 Concluding remarks
In this work we proposed a new soft tree model for survival analysis where the actual prediction for any input vector $\mathbf{x}$ is the corresponding survival function at the single leaf node obtained by routing the input vector from the root along the branches of higher probability. Such single leaf node predictions, which guarantee the conditional computation property, bring important benefits such as computational efficiency, modeling flexibility and interpretability. The proposed nonlinear optimization formulation is amenable to decomposition, and the node-based decomposition training algorithm with reassignment heuristic, originally proposed in (Consolo et al., 2025) for regression task, was adapted to the survival problem.
[…]
The above-mentioned results show that our SST model, whose leaf nodes correspond to clusters of distinct survival functions associated to different data points, the training formulation (2) and our NODEC-DR-SST algorithm offer a viable flexible method for survival analysis allowing both flexibility in the choice of survival functions and interpretability. SSTs, which consider multivariate soft splits, allow the use of any smooth survival function estimated via maximum likelihood. This is in contrast with the three recent survival tree approaches, based on MIO or dynamic programming, which only consider deterministic univariate splits (with further strong assumptions in Bertsimas et al. (2022); Huisman et al. (2024)) and survival functions that are restricted to specific non-parametric estimator (e.g., Kaplan-Meier).
As future work, it would be interesting to enhance the interpretability of SSTs by inducing sparsity to identify the most relevant features for prediction, as done in soft classification and regression trees (see, e.g., Blanquero et al. (2020); Amaldi et al. (2023); Blanquero et al. (2022)) via regularization or via a two-stage procedure (e.g., Carrizosa et al. (2023)). Furthermore, since SSTs trained via the proposed decomposition algorithm partitions the dataset into clusters, and each one of these clusters is associated to a single leaf node (with the same type of survival model), sparsity could also help to identify the relevant features that determine these partitions.
Another interesting direction would be to extend SSTs to handle more complex data, such as functional features, previously addressed for soft regression trees in Blanquero et al. (2023), spatial data Ripley (2005), or unstructured information such as text, thus enhancing their applicability.
@@ A.1 Datasets details
  • Aids: AIDS Clinical Trial (Lemeshow and May, 2008). The event refers to the onset of AIDS.
  • Aids__\__death: AIDS Clinical Trial (Lemeshow and May, 2008). The event refers to the death.
  • Aids2: Patients diagnosed with AIDS in Australia before 1 July 1991 (Ripley and Solomon, 1994). The event refers to death.
  • Churn: The task involves predicting when customers of a software-as-a-service (SaaS) company are likely to cancel their monthly subscription. The event refers to the end of the subscription. This dataset is from pysurvival (Fotso et al., 2019).
  • Credit: Lenders aim to predict when a borrower will repay their loan. The event refers to loan repayment. This dataset comes from PySurvival (Fotso et al., 2019) and has been adapted from the UCI Machine Learning Repository.
  • Dialysis: (SÁ Carvalho et al., 2003) . the survival of dialysis patients to assess the quality of renal replacement therapy at dialysis centers in Rio de Janeiro (SÁ Carvalho et al., 2003). The event refers to death. This dataset is from (Drysdale, 2022).
  • Employee: The task involves predicting when an employee will resign. The event refers to an employee’s departure. This dataset is from pysurvival (Fotso et al., 2019).
  • Flchain: Nonclonal serum immunoglobulin free light chains are used to predict overall survival in the population (Dispenzieri et al., 2012). The event refers to death.
  • Framingham: Survival after the onset of congestive heart failure in Framingham Heart Study (Mahmood et al., 2014). The event refers to the onset of heart failure. This dataset is from (Drysdale, 2022).
  • Gbsg2: The German Breast Cancer Study Group 2 (Schumacher et al., 1994) focuses on recurrence-free survival. The event refers to cancer recurrence-free survival.
  • Maintenance: The task involves predicting when equipment failure will occur. The event refers to machine failure. This dataset is from PySurvival (Fotso et al., 2019).
  • Uissurv: The UMASS AIDS Research Unit IMPACT Study (Lemeshow and May, 2008) focuses on the drug use. The event refers to the resumption of drug use.
  • Undempdur: The impact of changes in unemployment insurance disregards on job search behavior (McCall, 1996). This dataset is from (Drysdale, 2022). The event refers to reemployment.
  • Unemployement: Unemployment duration and job search theory, using revised Current Population Survey data from September–December 1993. This dataset is from (Drysdale, 2022). The event refers to reemployment. (Romeo, 1999).
  • Veterans: Veterans’ Administration Lung Cancer Trial (Kalbfleisch and Prentice, 2002). The event refers to death.
  • Whas500: Worcester Heart Attack Study (Lemeshow and May, 2008). The event refers to death.

### v3

@@ 5.3 Comparison with three benchmark survival tree methods
In the numerical experiments, we consider SSTs with parametric or spline-based semiparametric models for the survival functions. SSTs involving the parametric Exponential, Weibull, and LogLogistic distributions are denoted as Exp, W, and Llog, respectively. For the semiparametric survival function models, SSTs with the PO and PH models are referred to as PO and PH, respectively. To show the impact of an appropriate initial solution on the calibration and discrimination measures, we also report the results for the Llog parametric distribution when NODEC-DR-SST is combined with the K-means-based initialization procedure proposed in (consolo2025), referred to as Llog-init.
@@
Tables 3 and 4 report the average testing $C_{H}$ and IBS for each method. The arithmetic averages over all datasets appear at the bottom of each table. Within each group of approaches, namely parametric SST, semiparametric spline-based SST, and benchmark models for each dataset, the best model is highlighted in bold.
Concerning the $C_{H}$ discrimination measure in Table 3, for which higher values correspond to better solutions, Llog and Llog-init (with the clustering-based initialization) achieve the best average testing accuracy among the SSTs with parametric or semiparametric models at leaf nodes, and perform better than the three benchmarking methods. In particular, the average testing values for Llog, CTree (the best benchmarking method), and both the comparable semiparametric models PO and PH are, $C_{H}=0.744$, $C_{H}=0.711$, and $C_{H}=0.735$ and $C_{H}=0.736$. Overall, Llog outperforms CTree on 14 out of the 15 datasets, while PO and PH on 13 of them. By applying the Wilcoxon test between Llog and CTree, a p-value of $6.6e^{-4}$ is obtained. When comparing PO and PH with CTree, the resulting p-values are $7.3e^{-3}$ and $5.7e^{-3}$, respectively. These results confirm, at the $1\%$ significance level, the improvement of SSTs over the best benchmark model. The Employee dataset is the only one for which SkSurv, CTree, and RPART yield better $C_{H}$ values than the SST alternatives for all the four performance measures. Note that Llog-init, which includes an appropriate initialization procedure, leads to a one percent improvement with respect to Llog for the three discrimination measures, see B. The superiority of Llog-init in terms of accuracy is further supported by the p-values below $1\%$ obtained in the comparison with CTree for $C_{H}$, $C_{U}$, and CD-AUC, which are $3.79e^{-3}$, $5.3e^{-3}$, and $5.29e^{-1}$, respectively.
@@
Table 3: Testing accuracy in terms of $C_{H}$ discriminant measure (the higher the better). The comparison includes SSTs of depth $D=2$ with parametric distributions (Exp, W, Llog) and spline-based semiparametric models (PO and PH), as well as the three benchmarking survival tree methods (SkSurv, CTree, and RPART) with depth $D=5$. Llog-init refers to Llog with the clustering-based initialization procedure. Boldface indicates the best model within each group.
As to the calibration testing measure IBS in Table 4, for which lower values correspond to better solutions, the spline-based semiparametric SSTs, PO and PH, outperform the benchmarking methods SkSurv, CTree, and RPART, as well as the parametric SSTs. Among the benchmarking methods, CTree achieves the lowest IBS ($0.11$), comparable to Llog ($0.111$), but it turns out to be less accurate than PO and PH, which yield $\text{IBS}=0.074$ and $\text{IBS}=0.075$, respectively. Indeed, the Wilcoxon test comparing PO with CTree yields a p-value of $6.22e^{-3}$, which is below the $1\%$ significance level. Notice that the K-means-based initialization procedure further enhances the performance of Llog, achieving an average IBS of $0.107$ compared to $0.111$ with no initialization.
@@
Table 4: Testing accuracy in terms of calibration measure IBS (the lower the better). The comparison includes SSTs of depth $D=2$ with parametric distributions (Exp, W, Llog) and spline-based semiparametric models (PO and PH), as well as the three benchmarking survival tree methods (SkSurv, CTree, and RPART) with depth $D=5$. Llog-init refers to Llog with the clustering-based initialization procedure. Boldface indicates the best model within each group.
@@
Finally, to assess the computational efficiency of the NODEC-DR-SST algorithm, we evaluate the CPU times of SSTs with both parametric and non-parametric models for 15 datasets. Table 5 reports the mean computational time in seconds for each dataset over 100 runs of every SST model, with the corresponding median shown in brackets. The last row indicates the averages of the mean and of the median over all datasets. As expected, SSTs with parametric models require less computational time than SSTs with spline-based semiparametric ones. For all parametric models, computational times are below 88 seconds, except for the Dialysis and Employee datasets, whose mean and median reach up to 150 and 140 seconds, respectively. Specifically, for 11 datasets, both the mean and median computational times are consistently below 42 seconds. Overall, when averaged across all datasets, SSTs with parametric models do not require more than 35 seconds. For SSTs with spline-based models, the larger number of variables results in longer computational times. With the exception of the Aids2, Dialysis, Flchain, and Framingham datasets, which require computational times above 200 seconds, all other datasets need less than 57 seconds. For the Aids2 and Flchain datasets, SSTs with PO and PH models yield a median which is substantially lower than the mean. For instance, in the Aids2 dataset, PO achieves a median of $25$ seconds compared to an average of $234$ seconds, suggesting that only a small number of starting solutions result in longer computational times. The dataset with the highest computational burden is Flchain, for which PO and PH require on average $4844.13$ and $9148.55$ seconds, respectively. However, the median values are lower than the mean ones, and in the case of PO the median is more than an order of magnitude smaller. This indicates that the choice of the starting solution has a strong impact on the computation time. On average, PO requires smaller computational times than PH, with average means of $392.54$ and $711.01$ seconds, respectively, across all datasets. Notice that the averages are substantially affected by the computational times for Flchain. When Flchain is excluded, the average means for PO and PH are $74.57$ and $108.33$ seconds, respectively.
Dataset Llog Llog-init Exp W PO PH
Aids 4.86 (4.13) 14.78 (14.09) 4.20 (4.12) 6.29 (5.57) 23.92 (24.30) 34.82 (31.47)
Aids_death 5.25 (4.02) 12.46 (11.51) 4.72 (4.32) 5.43 (4.77) 28.82 (23.38) 29.69 (20.42)
Aids2 10.56 (8.52) 15.93 (16.11) 12.83 (11.53) 12.41 (10.97) 234.03 (24.97) 223.30 (18.90)
Churn 17.08 (18.42) 27.12 (26.94) 22.57 (24.31) 24.39 (24.39) 22.32 (23.06) 56.72 (52.33)
Credit 20.42 (21.22) 26.54 (26.99) 36.31 (36.39) 24.20 (24.65) 25.95 (25.66) 34.52 (33.03)
Dialysis 84.21 (49.43) 149.29 (166.67) 138.93 (126.24) 83.91 (52.65) 402.18 (359.56) 725.97 (699.34)
Employee 46.19 (43.56) 108.16 (117.98) 49.12 (45.85) 53.29 (49.32) 74.78 (70.86) 144.92 (139.37)
Flchain 68.87 (71.38) 87.80 (86.05) 87.84 (85.29) 76.21 (80.27) 4844.13 (209.06) 9148.55 (8688.07)
Framingham 18.80 (14.71) 32.78 (29.56) 42.64 (42.81) 24.01 (20.85) 192.26 (190.47) 212.38 (222.06)
Gbsg2 3.22 (3.14) 5.36 (5.64) 3.28 (3.41) 3.68 (3.56) 6.38 (6.04) 9.31 (8.68)
Maintenance 6.70 (6.97) 7.46 (7.57) 2.80 (2.66) 6.56 (6.55) 6.60 (6.58) 9.51 (9.48)
Uissurv 5.23 (5.47) 4.69 (4.28) 4.88 (4.89) 5.06 (5.07) 4.94 (4.77) 7.74 (7.74)
Unempdur 9.32 (8.80) 9.47 (8.99) 9.53 (8.57) 9.60 (8.92) 15.15 (15.53) 18.42 (19.12)
Veterans 1.01 (1.00) 2.61 (2.65) 0.98 (0.99) 1.04 (1.02) 1.29 (1.17) 2.28 (1.34)
Whas500 3.27 (3.26) 5.85 (5.95) 3.91 (3.91) 3.76 (3.78) 5.32 (4.87) 6.99 (6.81)
Average 20.33 (17.60) 34.02 (35.40) 28.30 (27.02) 22.66 (20.16) 392.54 (66.02) 711.01 (663.88)
Table 5: For each dataset, the mean computational time in seconds obtained by the SSTs with the specified survival functions model over 100 runs is reported, while the median is shown in brackets. The last row reports the average of the mean and of the median across the 15 datasets.
@@ 5.4 Enhancing interpretability
The aim of this section is to demonstrate how the SST model enhances interpretability by analyzing the clusters of survival functions corresponding to each leaf node. Unlike other single-tree methods, such as greedy algorithms (hothorn2015ctree) or more recent approaches employing advanced optimization techniques (bertsimas2022optimal; huisman2024optimal; zhang2024optimal), the SST is sufficiently flexible to define distinct survival functions for each data point within a leaf node. This approach enables the exploration of differences both across and within leaf node clusters. For the sake of conciseness, the analysis focuses on the $9$ dataset, with results for other $6$ datasets presented in the E.
@@ 5.5 Fairness
Formulation (12) is optimized using the node-based decomposition training algorithm on a socio-economic dataset where a sensitive group is identifiable. Since the datasets discussed in Section 5.1 involve applications where it is not natural to identify a sensitive group, this section focuses on the Unemployment dataset from romeo1999conducting. The Unemployment dataset consists of a four-month panel of revised Current Population Survey data from September 1993, where each observation corresponds to an unemployed individual searching for a job. Each row includes binary variables indicating the reason for unemployment (e.g., whether the individual is a job loser, job leaver, or labor force reentrant), the search method employed, and their demographic status. The dataset consists of $452$ data points, $7$ features, with a censoring level of $56.6\%$. Gender is selected as the sensitive feature for group SS [^6] This feature was not considered for SST training and hence, it can be used to predict survival functions of forthcoming individuals without even knowing this feature.
@@
For the sake of illustration, we consider SST with the Llog parametric distribution and a single fold of the 5-fold cross-validation. We also perform a single run without applying the clustering-based initialization procedure. It is important to highlight that any alternative parametric or semiparametric choice of survival function for the leaf nodes is feasible. The decision to present a single run is motivated solely by the aim of clearly illustrating the distribution of clusters across the 4 leaf nodes and its trend as the hyperparameter $\rho$, associated with the fairness penalty, increases. The $\rho$ hyperparameter starts from 0 and reaches $\frac{20}{N_{M}N_{F}}$ in five steps, defined by $\frac{1}{N_{M}N_{F}}[0,1,5,10,15,20]$, where $N_{M}$ and $N_{F}$ correspond to the number of males and females in the training dataset.
@@ 6 Concluding remarks
In this work we proposed a new soft tree model for survival analysis where the actual prediction for any input vector $\mathbf{x}$ is the corresponding survival function at the single leaf node obtained by routing the input vector from the root along the branches of higher probability. Such single leaf node predictions, which guarantee the conditional computation property, bring important benefits such as computational efficiency, modeling flexibility and interpretability. The proposed nonlinear optimization formulation is amenable to decomposition, and the node-based decomposition training algorithm with reassignment heuristic, originally proposed in (consolo2025) for regression task, was adapted to the survival problem.
The numerical experiments conducted on 15 well-known datasets demonstrate that our SSTs, using parametric and spline-based semiparametric survival functions and trained with the node-based decomposition Nodec-DR-SST algorithm, outperform the benchmark survival trees in terms of discrimination measures ($C_{H},\,C_{U},\,\text{CD-AUC}$) as well as calibration (IBS) one. Moreover, as illustrated on a real-world dataset related to unemployment, SSTs can be extended to take into account group fairness.
The above-mentioned results show that our SST model, whose leaf nodes correspond to clusters of distinct survival functions associated to different data points, the training formulation (7) and our NODEC-DR-SST algorithm offer a viable flexible method for survival analysis allowing both flexibility in the choice of survival functions and interpretability. SSTs, which consider multivariate soft splits, allow the use of any smooth survival function estimated via maximum likelihood. This is in contrast with the three recent survival tree approaches, based on MILP or dynamic programming, which only consider deterministic univariate splits (with further strong assumptions in bertsimas2022optimal; huisman2024optimal) and survival functions that are restricted to specific non-parametric estimator (e.g., Kaplan-Meier).
As future work, it would be interesting to enhance the interpretability of SSTs by inducing sparsity to identify the most relevant features for prediction, as done in soft classification and regression trees (see, e.g., blanquero2020sparsity; amaldi2021multivariate; blanquero2022sparse) via regularization or via a two-stage procedure (e.g., CARRIZOSA2023106180). Furthermore, since SSTs trained via the proposed decomposition algorithm partitions the dataset into clusters, and each one of these clusters is associated to a single leaf node (with the same type of survival model), sparsity could also help to identify the relevant features that determine these partitions.
Another interesting direction would be to extend SSTs to handle more complex data, such as functional features, previously addressed for soft regression trees in BLANQUERO2023106152, spatial data ripley2005spatial, or unstructured information such as text, thus enhancing their applicability.
Acknowledgment
This work was partially supported by project PID2022-137818OB-I00 (Ministerio de Ciencia e Innovación, Spain), with European Regional Development Fund (ERDF).
@@ A.1 Datasets details
  • Aids: AIDS Clinical Trial (lemeshow2008applied). The event refers to the onset of AIDS.
  • Aids_\_death: AIDS Clinical Trial (lemeshow2008applied). The event refers to the death.
  • Aids2: Patients diagnosed with AIDS in Australia before 1 July 1991 (ripley1994note). The event refers to death.
  • Churn: The task involves predicting when customers of a software-as-a-service (SaaS) company are likely to cancel their monthly subscription. The event refers to the end of the subscription. This dataset is from pysurvival (fotso2019pysurvival).
  • Credit: Lenders aim to predict when a borrower will repay their loan. The event refers to loan repayment. This dataset comes from PySurvival (fotso2019pysurvival) and has been adapted from the UCI Machine Learning Repository.
  • Dialysis: (sa2003survival) . the survival of dialysis patients to assess the quality of renal replacement therapy at dialysis centers in Rio de Janeiro (sa2003survival). The event refers to death. This dataset is from (drysdale2022).
  • Employee: The task involves predicting when an employee will resign. The event refers to an employee’s departure. This dataset is from pysurvival (fotso2019pysurvival).
  • Flchain: Nonclonal serum immunoglobulin free light chains are used to predict overall survival in the population (dispenzieri2012use). The event refers to death.
  • Framingham: Survival after the onset of congestive heart failure in Framingham Heart Study (mahmood2014framingham). The event refers to the onset of heart failure. This dataset is from (drysdale2022).
  • Gbsg2: The German Breast Cancer Study Group 2 (schumacher1994randomized) focuses on recurrence-free survival. The event refers to cancer recurrence-free survival.
  • Maintenance: The task involves predicting when equipment failure will occur. The event refers to machine failure. This dataset is from PySurvival (fotso2019pysurvival).
  • Uissurv: The UMASS AIDS Research Unit IMPACT Study (lemeshow2008applied) focuses on the drug use. The event refers to the resumption of drug use.
  • Undempdur: The impact of changes in unemployment insurance disregards on job search behavior (mccall1996unemployment). This dataset is from (drysdale2022). The event refers to reemployment.
  • Unemployement: Unemployment duration and job search theory, using revised Current Population Survey data from September–December 1993. This dataset is from (drysdale2022). The event refers to reemployment. (romeo1999conducting).
  • Veterans: Veterans’ Administration Lung Cancer Trial (kalbfleisch2002statistical). The event refers to death.
  • Whas500: Worcester Heart Attack Study (lemeshow2008applied). The event refers to death.